# Detecting and Interpreting NSFW Prompts in Text-to-Image Models through Uncovering Harmful Semantics

### v2

@@ 4.2 NSFW Prompts Detection
The larger the $Score(p)$, the more likely the current prompt contains NSFW semantics. By default, we use $Score(p)=0$ as the classification boundary. However, in practice, this threshold may be subject to a slight shift to account for the imbalanced data distribution. We discuss this in further detail in the Appendix A.3.
Recognizing that even powerful detection methods can be circumvented, we propose a red teaming framework to enhance defensive robustness. In this process, adversarial examples that successfully bypass the defense are collected and incorporated into the training set as a form of data augmentation. However, since red teaming often yields a limited number of successful samples, we assign higher weights to these instances during training to amplify their influence on the model. Experiments have shown that optimizing the NSFW feature through data augmentation can effectively reduce the success rate of adaptive attacks.
4.3 NSFW Prompts Interpretation
We further interpret NSFW prompts through a two-module framework. For the text modality, we identify the tokens that contribute most strongly to the NSFW semantics. For the image modality, we progressively attenuate the NSFW semantics at the hidden state level, revealing how the generated image gradually transitions to a benign state.
4.3.1 Text-Module Interpretation
In the text module, we interpret NSFW prompts by identifying NSFW tokens. A straightforward idea is to represent a token’s NSFW semantic association by directly calculating the similarity between its hidden state and the NSFW feature:
@@ 4.3.1 Text-Module Interpretation
However, this approach is flawed because the hidden state inherently contains semantic information from all preceding tokens. Therefore, we need to disentangle the individual contribution of each specific token. Similar to Equation 2, each token’s hidden state $c_{p_{i}}^{l,h}$ can be represented as a combination of $z_{j}^{l-1}$, with the corresponding attention weight $\alpha_{j,i}$ indicating the contribution of $z_{j}^{l-1}$.
$c_{p_{i}}^{l,h}=\sum_{j=0}^{N}\alpha_{j,i}^{l,h}W^{l,h}z_{j}^{l-1},$ (6)
where $\alpha_{j,i}^{l,h}$ are attention weights from the $j$-th token to the $i$-th token. For each layer, we need to use the $\alpha$ from the preceding layers to approximate the contribution of each token. Since $\{\alpha_{j,i}^{l,h}\}$ directly forms the attention map $A^{l,h}$, we multiply the attention maps to approximate the contribution of each token to the positions in the current layer. In this way, the interpretative results for each layer can be represented as:
$E(p)_{i}^{l,h}=\sum_{j=0}^{N}\{\prod_{k=0}^{l}A^{k,h}\}_{j,i}\hat{E}(p)_{i}^{l,h}$ (7)
By averaging the results from all attention heads, we can obtain the final interpretation for each token. Table 1 demonstrates several examples of our interpretation. Our method accurately identifies sensitive words in regular samples and adversarial tokens in adversarial samples.
Figure 4: Image-based interpretation with Stable Diffusion v1.4.
4.3.2 Image-Module Interpretation
In the image domain, we gradually attenuate the NSFW semantics contained in the prompts and examine the generated images to observe how the images evolve as the NSFW semantics are progressively diminished. Since the conditional embedding in T2I models includes embeddings of all tokens, we need to eliminate NSFW semantics from all tokens. Folling Bhalla et al. (2024), we assume that the feature space of each token is identical to that of the EOS token. This allows us to compute the NSFW score for each token by projecting the intermediate embeddings onto the NSFW features. Subsequently, we can attenuate the NSFW semantics of all tokens as follows:
@@ 4.3.2 Image-Module Interpretation
By recombining $\hat{c}_{p_{i}}^{l,h}$ to calculate $Z^{l}$, we can derive the modified conditional embedding. By gradually increasing the value of $\beta$, we can observe the process of NSFW semantics being progressively eliminated from the images. Taking the prompt “A beautiful naked woman.” as an example, Figure 4 illustrates the images generated for various values of $\beta$. As $\beta$ increases, the images gradually transition from NSFW to benign while maintaining the basic semantics of the prompt. In the Appendix A.4, we provide interpretations for more NSFW prompts and offer a more detailed analysis.
5 Experiments
5.1 Experimental Settings
@@
Datasets. The datasets we use comprise three categories. Clean datasets consist of benign data that does not generate NSFW images, including MSCOCO (Lin et al., 2014). Regular NSFW datasets include manually generated prompts with explicit NSFW semantics, includeing 4chan Prompts (Qu et al., 2023), I2P (Schramowski et al., 2023), NSFW200 (Yang et al., 2024c), and NSFW-LAION. Adversarial datasets comprise algorithmically generated adversarial NSFW prompts. We utilize samples generated by three adversarial attack methods: SneakyPrompt (Yang et al., 2024c), MMA (Yang et al., 2024a), and Ring-A-Bell (Tsai et al., 2024).
@@
Baselines. We select eight detection methods as baselines, encompassing both commercial and open-source models. Commercial models are OpenAI Moderation (OpenAI, 2023; Markov et al., 2023), Azure AI Content Safety (Microsoft Azure, 2024), AWS Comprehend (Amazon Web Services, 2025), and Aliyun Text Moderation (Alibaba, 2025). Open-source moderators include two lightweight models, NSFW-text-classifier (michellejieli, 2022) and Detoxify (unitaryai, 2022), as well as two SOTA methods, Latent Guard (Liu et al., 2024) and GuardT2I (Yang et al., 2024b).
@@
OpenAI Moderation 0.2976 0.0010 0.8220 0.4578 0.8616 0.7960 0.4974 1288.43
Azure AI Content Safety 0.4761 0.0118 0.8590 0.6302 0.7331 0.7708 0.4313 922.67
AWS Comprehend 0.4702 0.0730 0.8118 0.5576 0.7143 0.6244 0.2980 286.42
Aliyun Text Moderation 0.1736 0.0023 0.7897 0.2941 0.5856 0.6720 0.1799 99.16
NSFW-text-classifier 0.7325 0.3534 0.6699 0.5466 0.7627 0.6823 0.2922 9.14
Detoxify 0.5432 0.1778 0.7465 0.5379 0.7226 0.6455 0.3340 24.82
Latent Guard 0.5021 0.1403 0.7625 0.5346 0.7579 0.5995 0.1690 167.90
GuardT2I 0.7102 0.0779 0.8686 0.7318 0.9160 0.8207 0.3492 352.3
@@
Table 2: The overall evaluation of HiddenGuard.
Metrics. The metrics we use include TPR, FPR, Accuracy, F1 Score, AUROC, AUPRC, and TPR@FPR 1%. We set the threshold at the value that achieves the maximum F1 score in the training set. Moreover, we evaluate the efficiency of HiddenGuard by measuring the average time per query.
Implementation Details. We deploy HiddenGuard on three text encoders: CLIP-ViT-L (CLIP-L) (Openai, 2024), CLIP-ViT-bigG (CLIP-G) (laion, 2024), and T5-v1.1-XXL (T5) (google, 2024). They cover the vast majority of current open-source text-to-image models, as detailed in Table 6. We randomly selected 5,000 prompts from over 35,000 prompts to serve as the training set use the remaining as the test set. For more detailed experimental settings, please refer to Appendix A.8.
@@
OpenAI Moderation 0.0244 0.0600 0.8200 0.6800 0.3147 0.7030 0.6311 0.6117
Azure AI Content Safety 0.1184 0.2162 0.9920 0.8300 0.8287 0.8176 0.8058 0.9126
AWS Comprehend 0.1462 0.2303 1.0000 0.7900 0.5909 0.8977 0.7699 0.8252
Aliyun Text Moderation 0.0627 0.0415 0.9200 0.1700 0.3392 0.1502 0 0.5437
NSFW-text-classifier 0.4930 0.5841 1.0000 0.9700 0.7972 0.9644 0.9029 0.9806
Detoxify 0.2180 0.3384 1.0000 0.8300 0.4248 0.9377 0.7282 0.8544
Latent Guard 0.2451 0.3581 0.9720 0.3600 0.5455 0.7753 0.2427 0.5922
GuardT2I 0.5926 0.6157 0.8640 0.8100 0.7902 0.8365 0.8835 1.0000
@@
Table 3: The accuracy of each dataset.
5.2 Experiments Results
Table 2 presents the overall evaluation results across all datasets, with the best performance for each metric highlighted in bold. As shown in the table, HiddenGuard consistently outperforms previous approaches across almost all metrics. It demonstrates high detection accuracy across the CLIP-L, CLIP-G, and T5 models, proving its applicability to various text encoders. Table 3 shows accuracy across datasets, with the first six columns covering regular NSFW datasets and the last three focusing on adversarial datasets. The classification threshold is set to the value that maximizes the F1 score on the training set. HiddenGuard maintains high accuracy across all datasets, whereas the performance of other methods is inconsistent.
Additionally, we assess each method’s efficiency by measuring the average time per query, presented in the last column of Table 2. HiddenGuard requires significantly less time than other methods, primarily because they often utilize large models for detection. In contrast, HiddenGuard only incorporates multiple matrix operations during the text encoder’s inference process.
Figure 5: ROC curves of all methods.
To comprehensively illustrate accuracy across a range of thresholds, we plot the ROC curves for HiddenGuard and the baseline methods, as shown in Figure 5. ROC curve for HiddenGuard is consistently positioned above other methods. This demonstrates that HiddenGuard maintains superior performance compared to the baselines across a wide spectrum of classification thresholds.
@@ 5.2 Experiments Results
Ablation study. We discuss the impact of using NSFW features from only a single layer of the text encoder. Figure 6 presents the results when using each layer of CLIP-L for detection. Even when using attention heads from a single layer, many layers still achieve high accuracy. Similar results were obtained for CLIP-G and T5, which are presented in detail in the Appendix A.12. Although using attention heads from a single layer can achieve high accuracy, we recommend using the original HiddenGuard method for the highest precision in detection.
Furthermore, we demonstrate the effectiveness of HiddenGuard in both fine-grained classification and few-shot learning scenarios, with detailed results provided in the Appendix A.9 and A.10.
5.3 Adaptive Attack

### v3

@@ III-D NSFW Prompts Detection
The larger the $Score(p)$, the more likely the current prompt contains NSFW semantics.
Theoretically, if $\text{Score}(p)>0$, the prompt $p$ contains NSFW semantics and should be classified as an NSFW prompt. Conversely, if $\text{Score}(p)<0$, the prompt should be benign. However, experiments show that while setting the threshold to zero allows AEIOU to achieve high accuracy, optimal classification performance requires a slight threshold adjustment. We hypothesize that this is due to the distribution of the training set not fully representing the actual distribution of NSFW prompts, introducing bias in the NSFW features derived during training. In our experiments, we determine the threshold by selecting the one that yields the highest F1 Score on the training set. For different text encoders, the final offset ranges from 1% to 3%.
The approach above treats NSFW as a single comprehensive category. Suppose there is a need to subdivide it further or to identify specific categories of NSFW prompts, such as sex or violence. In that case, we can categorize NSFW prompts in the training set based on labels. This enables the calculation of NSFW features for each subcategory, facilitating the determination of NSFW scores for each. If we also need to detect NSFW prompts across all categories, we can aggregate the NSFW scores from all subcategories and use their maximum value as the final NSFW score.
Figure 4: Image-based interpretation with Stable Diffusion v1.4. In order to mitigate potential impact on the reader, we follow the established convention of prior work [yang2024mma, ba2023surrogateprompt] and obscure the NSFW images using both blurring and masking techniques.
III-E NSFW Prompts Interpretation
After the detection process, we further interpret NSFW prompts through a two-module framework. First, we develop an interpretative method to identify the tokens that most strongly contribute to NSFW semantics. Second, we investigate the generation mechanism of NSFW semantics in conditional embeddings by gradually attenuating the NSFW features in the hidden states. Using the resulting embeddings to generate images, we gain insights into how NSFW semantics can be progressively eliminated.
III-E1 Text-Module Interpretation
In the text module, we interpret NSFW prompts by identifying NSFW tokens. Within each attention head, we assess the NSFW semantic association at any given position by computing the cosine similarity between the hidden state at that position and the head’s NSFW feature:
@@ III-E1 Text-Module Interpretation
As tokens pass through the attention model, their semantics interrelate and intertwine, with each position’s hidden states encapsulating information from all preceding tokens. Consequently, the association between the current position’s hidden states and NSFW semantics cannot be directly used to represent the token’s connection to NSFW semantics accurately.
Similar to Equation 2, each position’s hidden state $c_{p_{i}}^{l,h}$ can be represented as a combination of $z_{j}^{l-1}$, with the corresponding attention weight $\alpha_{j,i}$ indicating the contribution of $z_{j}^{l-1}$.
$c_{p_{i}}^{l,h}=\sum_{j=0}^{N}\alpha_{j,i}^{l,h}W_{VO}^{l,h}z_{j}^{l-1},$ (11)
where $\alpha_{j,i}^{l,h}$ are attention weights from the $j$-th token to the $i$-th token. For Layer 1, since it has only undergone a single operation as described in Equation 11, we can reconstruct the actual contribution of each token based on $\{\alpha_{j,i}^{1,h}\}$:
$E(p)_{i}^{1,h}=\sum_{j=0}^{N}\alpha_{j,i}^{1,h}\hat{E}(p)_{i}^{1,h}.$ (12)
For deeper layers, we need to use the $\alpha$ from the preceding layers to approximate the contribution of each token. Since $\{\alpha_{j,i}^{l,h}\}$ directly forms the attention map $A^{l,h}$, we can multiply the attention maps from each layer to approximate the contribution of each token to the positions in the current layer. In this way, the interpretative results for each layer can be represented as:
$E(p)_{i}^{l,h}=\sum_{j=0}^{N}\{\prod_{k=0}^{l}A^{k,h}\}_{j,i}\hat{E}(p)_{i}^{l,h}.$ (13)
By aggregating the interpretation results from all attention heads, we can obtain the final interpretation for each token:
$E(p)_{i}=\frac{\sum_{l=1}^{L}\sum_{h=1}^{H}E(p)_{i}^{l,h}}{L\cdot H}.$ (14)
Table I demonstrates several examples of our interpretation. Our method accurately identifies sensitive words in regular samples and adversarial tokens in adversarial samples. We will conduct further experiments to demonstrate the effectiveness of the interpretability method.
III-E2 Image-Module Interpretation
In the image domain, we gradually attenuate the NSFW semantics contained in the prompts and examine the generated images to observe how the images evolve as the NSFW semantics are progressively diminished. Since the conditional embedding in T2I models includes embeddings of all tokens, we need to eliminate NSFW semantics from all tokens. We assume that the feature space of each token is identical to that of the EOS token. This allows us to compute the NSFW score for each token by projecting the intermediate embeddings onto the NSFW features. Subsequently, we can attenuate the NSFW semantics of all tokens as follows:
@@ III-E2 Image-Module Interpretation
By recombining $\hat{c}_{p_{i}}^{l,h}$ to calculate $Z^{l}$, we can derive the modified conditional embedding. By gradually increasing the value of $\beta$ and using conditional embedding to generate images, we can observe the process of NSFW semantics being progressively eliminated from the images.
We use two prompts to demonstrate the results of our image-module interpretation: the sexual prompt, ”a beautiful naked woman,” and the violent prompt, ”a man with blood all over his face.” Figure 4 shows the images generated with different values of $\beta$. As the value of $\beta$ increases, the images generated from both prompts gradually transition from NSFW to harmless. During this process, we observed two interesting phenomena. First, as the NSFW semantics are progressively reduced, the model still attempts to preserve the original semantics of the prompt. For the sexual prompt, the woman gradually turns away and conceals sensitive areas with her hands. Ultimately, only the area above her shoulders is visible, rendering the image harmless. Throughout, the image never violates the prompt’s directive.
For the violent prompt, the blood on the man’s face gradually changes color to resemble oil stains, and the marks themselves progressively diminish. However, even at $\beta=1$, some dirty marks remain on the man’s face. These examples illustrate that the prompt’s semantics are highly editable within the hidden states. By applying our explanation method, we can progressively eliminate the NSFW semantic content while preserving the original meaning of the prompt to the greatest extent possible.
IV Experiments
IV-A Experimental Settings
IV-A1 Datasets
@@
The datasets we use comprise three categories. Clean datasets consist of benign data that does not generate NSFW images. Regular NSFW datasets include manually generated prompts with explicit NSFW semantics. Adversarial datasets comprise algorithmically generated adversarial NSFW prompts.
Figure 5: ROC curves of all methods.
Clean Dataset. Our experiments use the validation captions of MSCOCO [lin2014microsoft] as the clean dataset. MSCOCO is a cross-modal image-text dataset, a popular benchmark for training and evaluating T2I generation models. We remove all captions containing sensitive words to ensure the samples are benign. A total of 25,008 captions are retained.
Regular NSFW Dataset. We gather data from multiple sources to comprehensively represent various types of NSFW semantics. I2P [schramowski2023safe] contains 4,703 NSFW prompts sourced from real users on Lexica [lexica]. The categories include hate, harassment, violence, self-harm, sex, shocking content, and illegal activities. Specifically, we categorize the prompts into hard and soft based on their level of harmfulness. 4chan Prompts [qu2023unsafe] contain 500 NSFW prompts collected from 4chan [4chan]. They predominantly encompass discriminatory and derogatory statements. NSFW200 [yang2024sneakyprompt] involves 200 NSFW prompts related to sexual and bloody content. NSFW-LAION contains 1,143 NSFW prompts we sampled from LAION-COCO [schuhmann2022laion], predominantly focused on sexual content.
  • I2P [schramowski2023safe] contains 4,703 NSFW prompts sourced from real users on Lexica [lexica]. The categories include hate, harassment, violence, self-harm, sex, shocking content, and illegal activities. Specifically, we categorize the prompts into hard and soft based on their level of harmfulness.
  • 4chan Prompts [qu2023unsafe] contain 500 NSFW prompts collected from 4chan [4chan]. They predominantly encompass discriminatory and derogatory statements.
  • NSFW200 [yang2024sneakyprompt] involves 200 NSFW prompts related to sexual and bloody content.
  • NSFW-LAION. We sample 1143 NSFW prompts from LAION-COCO [schuhmann2022laion] to enrich datasets. These prompts mainly focus on sexual content.
Adversarial Dataset. Adversarial datasets include adversarial samples obtained through three open-source adversarial attack methods: MMA [yang2024mma], SneakyPrompt [yang2024sneakyprompt], and Ring-A-Bell [tsai2023ring]. For MMA, we utilize the 1,000 successful adversarial samples provided by the authors. For the other two methods, we generate samples with the open-source algorithms, resulting in approximately 200 samples for each.
IV-A2 Baselines
@@
We select eight detection methods as baselines, encompassing both prompt-based and embedding-based approaches. These include four commercial models, two open-source models, and two SOTA methods. The commercial models are OpenAI Moderation [openai, markov2023holistic], Azure AI Content Safety [Azure], AWS Comprehend [AWS], and Aliyun Text Moderation [Ali]. These systems primarily rely on large transformer-based architectures to identify potential toxic content in prompts. The open-source moderators, NSFW-text-classifier [NSFWtext] and Detoxify [Detoxify], leverage lightweight models to detect NSFW text content, offering faster inference speeds. Latent Guard [liu2024latent] and GuardT2I [yang2024guardt2i] are embedding-based methods that represent the current SOTA in detecting NSFW prompts for T2I models.
TABLE II: The overall evaluation of AEIOU.
@@
OpenAI Moderation [openai] 0.2976 0.0010 0.8220 0.4578 0.8616 0.7960 0.4974 1288.43
Azure AI Content Safety [Azure] 0.4761 0.0118 0.8590 0.6302 0.7331 0.7708 0.4313 922.67
AWS Comprehend [AWS] 0.4702 0.0730 0.8118 0.5576 0.7143 0.6244 0.2980 286.42
Aliyun Text Moderation [Ali] 0.1736 0.0023 0.7897 0.2941 0.5856 0.6720 0.1799 99.16
NSFW-text-classifier [NSFWtext] 0.7325 0.3534 0.6699 0.5466 0.7627 0.6823 0.2922 9.14
Detoxify [Detoxify] 0.5432 0.1778 0.7465 0.5379 0.7226 0.6455 0.3340 24.82
Latent Guard [liu2024latent] 0.5021 0.1403 0.7625 0.5346 0.7579 0.5995 0.1690 167.90
GuardT2I [yang2024guardt2i] 0.7102 0.0779 0.8686 0.7318 0.9160 0.8207 0.3492 352.3
@@
[…]
IV-A3 Metrics
We evaluate our model using a standard suite of binary classification metrics, including the True Positive Rate (TPR), the False Positive Rate (FPR), overall Accuracy, and the F1 Score [powers2011evaluation]. For a threshold-agnostic analysis, we report the Area Under the ROC Curve (AUROC) for overall discriminative power and the Area Under the Precision-Recall Curve (AUPRC), which is more informative for skewed class distributions [fawcett2006introduction]. To assess performance under strict operational constraints, we also measure TPR at a low FPR of 1%. The decision threshold for these point-based metrics is determined by optimizing the F1 score on the training set. Finally, we assess the runtime efficiency of our method, AEIOU, by reporting its average inference time per query.
IV-A4 Implementation Details
We deploy AEIOU on three commonly used text encoders: CLIP-ViT-L (CLIP-L) [clipl], CLIP-ViT-bigG (CLIP-G) [clipg], and T5-v1.1-XXL (T5) [t5]. CLIP-L is the most widely used text encoder and serves as the foundation for most safety-related research. CLIP-G and T5 are larger models; many recent models [flux, esser2024scaling] employ them as text encoders.
We design two variant detectors to evaluate our method’s effectiveness in different scenarios. First, when the defender is unaware of the attack method, we employ a model trained solely on clean data and regular NSFW data to test AEIOU’s generalization capability. Second, when the defender has sufficient data labeled with specific categories, they can train on multiple specific concepts to enhance the model’s specificity. We deploy the CLIP-L-based detector in both scenarios.
TABLE III: The accuracy of each dataset.
@@ IV-A4 Implementation Details
OpenAI Moderation [openai] 0.0244 0.0600 0.8200 0.6800 0.3147 0.7030 0.6311 0.6117
Azure AI Content Safety [Azure] 0.1184 0.2162 0.9920 0.8300 0.8287 0.8176 0.8058 0.9126
AWS Comprehend [AWS] 0.1462 0.2303 1.0000 0.7900 0.5909 0.8977 0.7699 0.8252
Aliyun Text Moderation [Ali] 0.0627 0.0415 0.9200 0.1700 0.3392 0.1502 0 0.5437
NSFW-text-classifier [NSFWtext] 0.4930 0.5841 1.0000 0.9700 0.7972 0.9644 0.9029 0.9806
Detoxify [Detoxify] 0.2180 0.3384 1.0000 0.8300 0.4248 0.9377 0.7282 0.8544
Latent Guard [liu2024latent] 0.2451 0.3581 0.9720 0.3600 0.5455 0.7753 0.2427 0.5922
GuardT2I [yang2024guardt2i] 0.5926 0.6157 0.8640 0.8100 0.7902 0.8365 0.8835 1.0000
@@
[…]
IV-B Overall Evaluation
We first conduct an overall evaluation of the effectiveness of AEIOU. We deploy it across three different text encoders and trained models using three distinct settings. It is then compared against eight baselines on one benign and eight NSFW datasets. We randomly selected 2,000 prompts from a total of over 35,000 prompts to serve as the training set use the remaining as the test set.
Table II presents the overall evaluation results across all datasets, with the best performance for each metric highlighted in bold. As shown in the table, AEIOU consistently outperforms previous classification approaches across almost all metrics. It demonstrates high detection accuracy across the CLIP-L, CLIP-G, and T5 models, proving its applicability to various text encoders.
[…]
Additionally, we assess each method’s efficiency by measuring the average time per query, presented in the last column of Table II. AEIOU requires significantly less time than other methods, primarily because they often utilize large models for detection. In contrast, AEIOU only incorporates multiple matrix operations during the text encoder’s inference process. On the smallest CLIP-L model, AEIOU’s efficiency improves at least tenfold compared to other models. Even on larger models like CLIP-G and T5, AEIOU’s efficiency surpasses that of all other models.
To comprehensively illustrate accuracy across a range of thresholds, we plot the ROC curves for AEIOU and the baseline methods, as shown in Figure 5. ROC curve for AEIOU is consistently positioned above other methods. This demonstrates that AEIOU maintains superior performance compared to the baselines across a wide spectrum of classification thresholds.
IV-C Generalization to Unknown Attacks
[…]
TABLE IV: The evaluation of different categories.
AEIOU 0.9959 0.9863 0.9763 0.9921 0.9790 0.9733 0.9340
[…]
TABLE V: The impact of training data size.
Training Data Size TPR FPR ACC F1 Score AUROC AUPRC TPR@FPR 1%
10 0.9171 0.0278 0.9583 0.9174 0.9875 0.9734 0.8495
50 0.9356 0.0164 0.9714 0.9430 0.9940 0.9861 0.9157
100 0.9605 0.0180 0.9766 0.9538 0.9962 0.9906 0.9351
500 0.9824 0.0091 0.9888 0.9778 0.9981 0.9974 0.9829
1000 0.9833 0.0096 0.9887 0.9776 0.9990 0.9975 0.9833
Figure 6: The impact of training data size.
IV-D Multi-Categories Classifier
[…]
IV-E The Impact of Training Data Size
@@
TABLE VI: The evaluation of adaptive attack.